# Probing the cosmic sterile-neutrino background with IceCube

### v1

@@ III Searching for dips in spectrum
III.1 Forecast for IceCube-Gen2
IceCube-Gen2 is the planned upgrade of the IceCube neutrino observatory, which is expected to have $\sim$10 times larger exposure to TeV–PeV neutrinos Aartsen et al. (2021b). To estimate the event rate in IceCube-Gen2, we scale the weights according to the increased effective area. The effective area for cascades in IceCube-Gen2 is given in Ref. Aartsen et al. (2021b). However, the effective area for tracks in IceCube-Gen2 is not publicly available at the time of writing this paper. In this work, we will assume that the effective area is the same for cascades and tracks. We have evaluated the event rates using 25 years of IceCube-Gen2 exposure for the best-fit parameters from the SPL analysis and for the best-fit point from this work. The results are presented in Figure 5. The local significance of the dips can be tested with such large exposures. It must be noted that there are other proposed neutrino telescopes that will also be sensitive to the absorption features. A combined analysis will improve the discovery prospects Fiorillo and Bustamante (2023).
@@ IV Summary and outlook
The IceCube neutrino observatory has played a crucial role in our understanding of astrophysical neutrinos in the last decade. It also provides an opportunity to study rare interactions over large distances. In this paper, we have taken a close look at the interaction of the astrophysical neutrinos with a hypothetical sterile neutrino background. The light sterile neutrino is well-motivated from short baseline experiments. However, there is a conflict with standard cosmology, and one of the resolutions is to allow self-interactions via a MeV-scale mediator. As a consequence, one expects the cosmic neutrino background to have a sterile neutrino component as well. The astrophysical neutrinos can interact with the sterile neutrino background, which leads to dips in the TeV–PeV neutrino energy spectrum. If the dips are significant, there is an observable impact on the HESE sample in IceCube.
We have improved on the previous analysis on multiple fronts. We have taken a realistic source distribution and included the effect of regeneration during the propagation. We also account for the recent observation of Glashow resonance in our analysis. To estimate the likelihood, we first compute the event rate and then compare with the observations. We find that the inclusion of these absorption features allows for a lower spectral index ($\gamma_{\texttt{astro}}=2.60^{+0.19}_{-0.16}$) when compared with the single-power law flux ($\gamma_{\texttt{astro}}=2.87^{+0.20}_{-0.19}$). Note that, due to the approximations in our analysis, our errors are underestimated. The lower spectral index is in better agreement with the Northern Tracks sample ($\gamma_{\texttt{astro}}=2.28^{+0.08}_{-0.09}$). However, a detailed study that accounts for the effect of these absorption features on the Northern Tracks sample is essential and left as future work. Currently, with only 7.5 years of HESE data, we find that there is a large overlap of parameter space, and one cannot statistically infer the existence of such dips. In the future, with a much larger neutrino telescope in the form of IceCube-Gen2 and perhaps other IceCube-scale experiments, the absorption features will be strongly tested.
@@ Acknowledgments
The authors would like to thank Austin Schneider for helping with the HESE Monte Carlo simulation and Ivan Esteban for helping with nuSIprop. The authors would also like to thank Subhendra Mohanty and Ivan Esteban for their comments and suggestions on the manuscript. PP thanks Ranjan Laha for useful discussion. BC is supported by an OPERA grant and seed grant NFSG/PIL/2023/P3821 from BITS Pilani. PP acknowledges the support from the National Post-Doctoral Fellowship by the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (SERB-PDF/2023/002356) and the IOE-IISc fellowship program.

### v2

@@ IV Summary and outlook
The IceCube neutrino observatory has played a crucial role in our understanding of astrophysical neutrinos in the last decade. It also provides an opportunity to study rare interactions over large distances. In this paper, we have taken a close look at the interaction of the astrophysical neutrinos with a hypothetical sterile neutrino background from the perspective of various IceCube datasets. The light sterile neutrino is well-motivated from short baseline experiments. However, there is a conflict with standard cosmology, and one of the resolutions is to allow self-interactions via a MeV-scale mediator. As a consequence, one expects the cosmic neutrino background to have a sterile neutrino component as well. The astrophysical neutrinos can interact with the sterile neutrino background, which leads to dips in the TeV–PeV neutrino energy spectrum. If the dips are significant, there is an observable impact on the starting events such as HESE and ESTES. The Northern Tracks are not significantly affected as there is a lack of one-to-one correspondence between observed event energy and the incident neutrino energy.
We have improved on the previous HESE analysis on multiple fronts as well. We have taken a realistic source distribution and included the effect of regeneration during the propagation. We also account for the recent observation of Glashow resonance in our analysis. To estimate the likelihood, we first compute the event rate and then compare with the observations. We find that the inclusion of these absorption features allows for a lower spectral index ($\gamma_{\texttt{astro}}=2.60^{+0.19}_{-0.16}$) when compared with the single-power law flux ($\gamma_{\texttt{astro}}=2.87^{+0.20}_{-0.19}$). Note that, due to the approximations in our analysis, our errors are underestimated. The lower spectral index is in better agreement with the Northern Tracks sample ($\gamma_{\texttt{astro}}=2.37^{+0.09}_{-0.09}$), which are largely insensitive to the presence of such narrow features in the neutrino spectrum. The ESTES sample can prove complementary probe of these dips and can provide information about the flavor structure, if any. Currently, with only 7.5 years of HESE data, we find that there is a large overlap of parameter space, and one cannot statistically infer the existence of such dips. In the future, with a much larger neutrino telescope in the form of IceCube-Gen2 and perhaps other IceCube-scale experiments, the absorption features will be strongly tested.
@@ Acknowledgments
The authors would like to thank Austin Schneider for helping with the HESE Monte Carlo simulation and Ivan Esteban for helping with nuSIprop. The authors would also like to thank Subhendra Mohanty and Ivan Esteban for their comments and suggestions on the manuscript. The authors thank Ranjan Laha for useful discussions. The authors acknowledge the anonymous referee who motivated a large body of work in this manuscript. BC is supported by an OPERA grant and seed grant NFSG/PIL/2023/P3821 from BITS Pilani. PP acknowledges the support from the National Post-Doctoral Fellowship by the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (SERB-PDF/2023/002356) and the IOE-IISc fellowship program. The authors acknowledge the use of computational resources provided by DIST-FIST Project No. SR/FST/PS-1/2017/30.
Appendix A
In order to estimate the effects of our approximations on the fit, we evaluate the 68% C.L. and 90% C.L. intervals for single-power law and compare them with the contours in Ref. Abbasi et al. (2021a). The results are shown in Figure 8. We only scan over the astrophysical flux parameters and keep other parameters fixed close to their best-fit values. We find that on expected lines, the best-fit point agrees very well, and the errors are underestimated. It is also seen that the impact of these approximations lead to a larger change in the normalization and the impact on the spectral index is small.
Figure 8: The SPL fit to HESE data Abbasi et al. (2021a) considering all nuisance parameters (see Table 1) is shown in grey. The inner (dark grey) region shows the 68 % C.L. interval and the outer (light grey) region shows the 90 % C.L. interval. The best-fit point is shown as a grey star. Our fit to HESE data in which the nuisance parameters are fixed to their nominal values is shown as red curves. Our best-fit point for HESE-only analysis is very close to the IceCube best-fit point and not shown here. The fit including HESE and Glashow Resonance (GR) is shown in blue. The blue star shows the best-fit point of our HESE+GR analysis.
The impact of including the Glashow Resonance event is also on expected lines. In order to get a reasonable expectation for the energy bin, one must have a harder spectrum, and larger values of spectral index are not preferable. As one can see in the Figure 8, the fits for smaller values of $\gamma_{\texttt{astro}}$ for both HESE and HESE+GR are similar, and differ for larger values.